OpenStudy (anonymous):

Someone please help teach me this. MEDAL AND FAN!!!! The expression below shows the population of squirrels in a forest after 3x years: f(x) = 30(1.03)3x Which of the following is an equivalent function? f(x) = 30(3.0900)x f(x) = 30(1.0927)x f(x) = 10(1.0927)x f(x) = 10(3.0900)x

OpenStudy (jdoe0001):

\(\bf f(x) = 30(1.03)3x\quad or\quad f(x) = 30(1.03)3^x\quad or\quad f(x) = 30(1.03)^{3x}?\)

OpenStudy (anonymous):

@jdoe0001 The last one

OpenStudy (jdoe0001):

\(\bf (a^n)^{m}\iff a^{n\cdot m} \\ \quad \\ \quad \\ f(x) = 30(1.03)^{{\color{blue}{ 3}}x}\implies f(x) = 30(1.03^{\color{blue}{ 3}})^{x}\implies ?\)

OpenStudy (anonymous):

@jdoe0001 Man, I have no idea how to do this. Are their certain steps I am supposed to follow? If so can i get an example

OpenStudy (jdoe0001):

well.... what would that exponentialized value give inside the parentheses?

OpenStudy (anonymous):

1.092727? @jdoe0001

OpenStudy (jdoe0001):

yeap.... thus \(\bf f(x) = 30(1.03)^{{\color{blue}{ 3}}x}\implies f(x) = 30(1.03^{\color{blue}{ 3}})^{x}\implies f(x)=30(1.092727)^x\)

OpenStudy (jdoe0001):

and you can round that up to only say 1.0927

OpenStudy (anonymous):

So we have f(x) = 30(1)^x right?

OpenStudy (anonymous):

@jdoe0001